Metal A has a heat capacity of 0.450 J/go C while metal B has a heat capacity of 1.250 J/go C. If both metals are supplied with 1000.0 J of energy, which one will heat up more? Explain by describing what specific heat capacity defines
1
Expert's answer
2020-06-04T10:29:55-0400

The specific heat capacity of a substance is the amount of energy required to raise the temperature of 1 g of a substance by 1 degree Celsius.

If we have 1 g of each of the metals A and B, and the amount of heat supplied for each of the metals is 1000.0J, then:

find the raise in temperature for each of the metals:

"q=cm\Delta T"

"\Delta T = \frac{q}{cm}"


"\Delta T (A) = \frac{1000. 0J}{0.450 \frac{J}{g\times ^\circ C}\times 1 g}=2222.22 ^\circ C"

"\Delta T(B) = \frac{1000.0 J}{1.250 \frac{J}{g\times ^\circ C}\times 1 g}=800 ^\circ C"

So, we can see that the smaller the heat capacity of a metal the more it will be heated when the equal amount of heat is supplied to both metals.